OpenStudy (anonymous):

Algebra 2 Help please!! I will give a medal. IN WHAT DIFFERENT WAYS CAN EXPONENTIAL FUNCTIONS BE IDENTIFIED?

OpenStudy (solomonzelman):

Do you have a multiply choice to the question, is this a question that you are supposed to address in your own words, or are you just trying to learn about exponential functions on your own?

OpenStudy (anonymous):

@SolomonZelman It is just a question that I have to answer. Can you please help me?

OpenStudy (solomonzelman):

I will assume you know what exponent is.

OpenStudy (anonymous):

Yes, I do @SolomonZelman

OpenStudy (solomonzelman):

So, for example: \(\color{#000000 }{ \displaystyle 2^1=2 }\) \(\color{#000000 }{ \displaystyle 2^2=2\cdot 2 }\) (2 times) \(\color{#000000 }{ \displaystyle 2^3=2\cdot 2\cdot 2 }\) (3 times) \(\color{#000000 }{ \displaystyle 2^4=2\cdot 2\cdot 2 \cdot 2}\) (4 times) \(\color{#000000 }{ \displaystyle 2^x=2~~~{....}~~~ 2 \cdot 2}\) (x times) So the more times you multiply times 2 our output will increase. And this is not just when you multiply times 2. Every time you multiply times a number that is greater than 1 (let's call it b), you get a greater result. correct?

OpenStudy (solomonzelman):

So, if you make a function, \(\color{#000000 }{ \displaystyle y=b^x}\) where b is a number greater than 1, than as x increases, the output (y) would also increase.

OpenStudy (anonymous):

what does the output mean ? @SolomonZelman

OpenStudy (solomonzelman):

output is the result you get for y

OpenStudy (solomonzelman):

input is the x value that you plug into the function

OpenStudy (solomonzelman):

And any function in a form of \(\color{#000000 }{ \displaystyle y=b^x}\) (where b is some number greater than 1) (or in a form of \(\color{#000000 }{ \displaystyle y=a(b)^x}\), when the function is multiplied times a scale factor "a") THEN, your output will increase when input increases.

OpenStudy (solomonzelman):

So functions like this \(\color{#000000 }{ \displaystyle y=b^x}\) \(\color{#000000 }{ \displaystyle y=a(b)^x}\) (where b is greater than 1) is called "exponential growth"

OpenStudy (solomonzelman):

When x is bigger, you multiply times b more times. And since be is greater than 1, multiplying times 1 will yield larger and larger results. This is why such functions are called exponential growth

OpenStudy (solomonzelman):

are you getting what I am saying? any confusions so far?

OpenStudy (anonymous):

I am getting it so far but I am re-reading it all. @SolomonZelman

OpenStudy (solomonzelman):

Ok, sure... tell me when you are done, and if you have any questions ask, if not I'll proceed

OpenStudy (anonymous):

So, exponential growth is one of the ways to identified exponential functions, right? Are there more ways?

OpenStudy (anonymous):

And If I graph exponential growth that will be a not straight line, right? @SolomonZelman

OpenStudy (solomonzelman):

No it won't although it will tend to be a vertical line as x grows, because its rate of change is increasing.

OpenStudy (solomonzelman):

it's slow will grow as large as you wish, and will "approach infinity"

OpenStudy (solomonzelman):

Which one is smaller \(\color{#000000 }{ \displaystyle 0.5^1 }\) or \(\color{#000000 }{ \displaystyle 0.5^4 }\) and tell me why?

OpenStudy (anonymous):

the one with the exponent of 4 is the bigger because 0.5 will get larger as I times it 4 times so 0.5^1 will be the smaller. Am I right? That's what my understanding from what you told me @SolomonZelman

OpenStudy (solomonzelman):

0.5 is same as 1/2

OpenStudy (solomonzelman):

when you multiply 1 times 1/2 you will get a result that is smaller, wouldn't you?

OpenStudy (solomonzelman):

What is smaller, \(\color{#000000 }{ \displaystyle 1\times (1/2) }\) or 1 ?

OpenStudy (anonymous):

1x(1/2) is smaller @SolomonZelman

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

So I will list the products in order, and you will tell me what happens to them.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 1\times (1/2) }\) \(\color{#000000 }{ \displaystyle 1\times (1/2)\times (1/2) }\) \(\color{#000000 }{ \displaystyle 1\times (1/2)\times (1/2) \times (1/2) }\) \(\color{#000000 }{ \displaystyle 1\times (1/2)\times (1/2) \times (1/2)\times (1/2) }\) \(\color{#000000 }{ \displaystyle 1\times (1/2)\times (1/2) \times (1/2) \times (1/2)\times (1/2) }\)

OpenStudy (solomonzelman):

the product is increasing?

OpenStudy (solomonzelman):

or is it decreasing?

OpenStudy (anonymous):

it is increasing @SolomonZelman

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

it is decreasing. tell me why is it decreasing

OpenStudy (anonymous):

oh its because of the one half

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Explain

OpenStudy (anonymous):

because it you times 0.5 to 0.5 the results that you will get will be half of the number and it will be the same for the rest @SolomonZelman

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

So, if you set, \(\color{#000000 }{ \displaystyle y=(0.5)^x}\) then as x grows, what happens to y? (explain)

OpenStudy (anonymous):

It decreases because you are multiplying the x in a half @SolomonZelman

OpenStudy (solomonzelman):

The y decreases, because the bigger x is, the more times you multiply times (1/2), and therefore the more the result becomes smaller.

OpenStudy (solomonzelman):

And in fact, (should be understandable), that \(\color{#000000 }{ \displaystyle y=b^x}\) (if b is positive but less than 1) then, when you multiply times this "b" the more you minimize your result. That means that the bigger the x, the more you multiply times b, and therefore the smaller tha y you have.

OpenStudy (solomonzelman):

So a function is a form \(\color{#000000 }{ \displaystyle y=b^x}\) (if: 0>b>1 -- positive b, but smaller than 1) then as x tends to larger values, the y tends to 0. And therefore this function is called "exponential decay"

OpenStudy (anonymous):

is the exponential decay also part of how to identify an exponential functions? @SolomonZelman

OpenStudy (solomonzelman):

when I wrote 0>b>1, I mean to write 0<b<1

OpenStudy (solomonzelman):

well, yes.

OpenStudy (solomonzelman):

There are two basically

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y=b^x}\) or \(\color{#000000 }{ \displaystyle y=a(b)^x}\) when b is greater than 1. (each time you multiply times b, the result increases, so these two functions are called "exponential growth") \(\color{#000000 }{ \displaystyle y=b^x}\) or \(\color{#000000 }{ \displaystyle y=a(b)^x}\) when b ispositive, but smaller than 1. (each time you multiply times b, the result decreases, so these two functions are called "exponential decay"")

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y=b^x}\) is same as \(\color{#000000 }{ \displaystyle y=1\cdot (b)^x}\) so you can write just the \(\color{#000000 }{ \displaystyle y=a(b)^x}\) for eac one. you don't need to list both

OpenStudy (anonymous):

@SolomonZelman Thank you so much!! It makes it easier for me to understand.